The data below was collected from random samples of high school students in various counties given a new standardized test in science. SCORES LOCATION Fresno County L.A. County Kern County Sacramento County Fail Low Pass Pass High Pass 30 60 70 150 15 40 50 80 150 400 90 250 75 80 30 50 13.What sort of scale of measurement was used to record the scores of the students(Nominal, ordinal, interval or ratio)? 14.What sort of variable is represented by the location designations? 15.What proportion of the entire sample of students scored a High Pass on the test? 16.Among those that failed the test, what proportion were from Sacramento County? 17.Among those from L.A., what proportion scored a High Pass on this test? 18.What proportion of the entire sample scored a Low Pass and were from Kern County?
Answer #1

Ans:

13)Scores are measured on Ordinal scale.

14)Categorical or qualitative variable(Nominal scale)

Fail Low pass Pass High pass Total
Fresno 30 60 150 75 315
L.A 70 150 400 80 700
Kern 15 40 90 30 175
Sacromento 50 80 250 50 430
Total 165 330 890 235 1620

15)

P(high pass)=235/1620=0.145

16)

P(Sacromento)=50/165=0.303

17)

P(high pass/L.A)=80/700=0.114

18)

P(low pass and kern)=40/1620=0.025